five years ago i was thrice as old as my son and ten years later i shall be thrice as old as my son. what shall be the present age of us?
Answer by stanbon(75887)

five years ago i was thrice as old as my son and ten years later i shall be thrice as old as my son. what shall be the present age of us?
-----
(m-5) = 3(s-5)
m+10 = 3(s+10)
----------------------
Rearrange:
m - 3s = -10
m - 3s = 20
---
That is a contradiction.
Please check the wording of your problem.
Cheers,
Stan H.